OpenStudy (anonymous):

Brenna had a gross income of $2158.22 during each pay period in 2010. If she got paid monthly, how much of her pay was deducted for FICA in 2010?

OpenStudy (anonymous):

A. $1335.16 B. $1723.09 C. $1921.15 D. $1981.25

OpenStudy (anonymous):

D

OpenStudy (anonymous):

^^ so its D ?

OpenStudy (tkhunny):

Not sure. What's the tax rate or schedule?

OpenStudy (anonymous):

Her annual gross income is 2158.22∗12=25898.64. So, you need to calculate 7.65% of that. Therefore, the answer is: 25898.647.65100=$1981.24596 Which is D

OpenStudy (tkhunny):

Not sure what number I was using, before. Let's try this. 2158.22*12 = 25898.64 1335.16/25898.64 = 0.05155 1723.09/25898.64 = 0.06653 1921.15/25898.64 = 0.07418 1981.25/25898.64 = 0.07650 Well, then there you go.